### v1

# Private Online Community Detection for Censored Block Models

Abstract
We study the private online change detection problem for dynamic communities, using a censored block model (CBM). Focusing on the notion of edge differential privacy (DP), we seek to understand the fundamental tradeoffs between the privacy budget, detection delay, and exact community recovery of community labels. We establish the theoretical lower bound on the delay in detecting changes privately and propose an algorithm capable of identifying changes in the community structure, while maintaining user privacy. Further, we provide theoretical guarantees for the effectiveness of our proposed method by showing necessary and sufficient conditions on change detection and exact recovery under edge DP. Simulation and real data examples are provided to validate the proposed method.
1 Introduction
Community detection in networks is a fundamental problem in the area of graph mining and unsupervised learning [14]. The main goal is to find partitions (or communities) in the graph that have dense connections within communities and sparse connections across communities. The Stochastic Block Model (SBM) is commonly used to model the graphical structure for networks [2, 17]. In SBM model, nodes are divided into communities, and the connection probability between nodes depends on whether they belong to the same communities or not. Connections between nodes are more probable within the same community than between communities. This difference in connection probabilities is central to the challenge of community detection. The study and improvement of community detection methods using the SBM framework has been a dynamic area of research, with many advances and findings detailed in comprehensive surveys such as the one by Abbe et al. [1].
In contemporary research on network community detection, a significant focus has been placed on static network structures. However, this focus on static networks does not completely capture the dynamic nature of real-world networks, where the arrangement of nodes and their links constantly changes. Despite some advances (e.g., [37, 3]) in applying the dynamic SBM to these scenarios, a comprehensive understanding of the conditions that allow for accurate community detection in such dynamic environments, particularly within the framework of the dynamic Censored Block Model (CBM), remains an open and critical area of research.
The study of CBM in dynamical networks has become increasingly relevant in network science, offering critical insights into the dynamics of complex systems [16, 25]. CBM models effectively represent real-world networks characterized by block structures and are particularly valuable in analyzing networks with incomplete or censored data. This feature is essential for understanding various network dynamics, such as information spread and epidemic outbreaks. CBM models are also instrumental in the development and testing of network science algorithms, especially for community detection and network reconstruction.
The information-theoretic conditions for recoverability are well understood in terms of the scaling of the CBM parameters $p$ and $\zeta\in(0,1/2)$, where $p$ represents the probability of an edge being observed, while $\zeta$ denotes the connection probability (these quantities will be defined more precisely in the sequel). In particular, in the dense regime (which is the focus of this paper), with $p=a\log(n)/n$ for some constant $a>0$, it is known that exact recovery under binary CBMs is possible if and only if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$ [18]. The condition above is considered to be a sharp threshold for recovery, which can also be interpreted as a phase transition. In the literature, efficient algorithms for recovering communities have been developed using spectral methods (e.g., [10, 9]) and semi-definite programming (SDP) [17]. Some of these results were proven to achieve the optimal threshold for exact recovery. In contrast, the sparse regime, defined by $p=a/n$, permits only partial recovery, meaning the estimator is positively correlated with the ground truth labels, provided that $a(1-2\zeta)^{2}>1$, as demonstrated in [21]. Furthermore, spectral algorithms have been shown to achieve this sharp threshold efficiently, as evidenced in [31].
In this work, we study the dynamic CBM models under privacy constraints, since sensitive information is frequently present in network data. As such, ensuring the privacy of individuals during data analysis is imperative. Differential Privacy (DP) [12] has become the recognized standard for delivering robust privacy assurances. DP guarantees that any single user’s participation in a dataset only minutely impacts the statistical outcomes of queries. In the literature [20], two main privacy notions have been proposed for data analysis over graphs: $(1)$ edge DP that aims to protect the individual relationships (edges) in a graph, using randomized algorithms, and $(2)$ node DP that focuses of the privacy of the nodes and their corresponding interactions (edges). Notably, DP algorithms have been tailored to specific network analysis problems, including counting stars, triangles, cuts, dense subgraphs, and communities or even generating synthetic graphs, have been introduced [27, 30, 19, 5].
Under edge DP constraints, we study the online change detection of community structures for CBM models. We propose a private online detection algorithm that performs window-based detection on privatized data. Sufficient conditions for exact recovery of the unknown community structure, under DP constraints, are established. Furthermore, we provide a theoretical guarantee for the detection performance of the proposed method and a new fundamental limit to private exact recovery for the community structure of the CBM.
Related Works
The problem of private change detection is an evolving field with significant applications. It has been studied under $(\epsilon,\delta)$ differential privacy constraints for multiple settings [8, 7]. In [4], the online change detection problem has been studied under the locally private condition. In [22], the authors extended the analysis of privatized networks from static to dynamic, focusing on sequences of networks with change-points. They explored the limits of consistently localizing change points under edge privacy constraints, uncovering a phase transition in the signal-to-noise ratio conditions, and developing polynomial-time algorithms. The research highlighted how privacy influenced the efficiency of change point localization, with a notable difference in the scaling of sparsity parameters compared to non-private scenarios. The algorithms were found to be near-optimal under edge Local Differential Privacy (LDP) constraints. This work underscored the complexities and challenges of preserving privacy in the dynamic analysis of network data.
In contrast, our paper broadens the scope of analysis by considering possible network changes, in particular, shifts in community memberships. This approach presents a more comprehensive and challenging problem, as it involves not only the detection of statistical changes but also the identification of structural and compositional alterations within the network dynamics. Our methodology, therefore, addresses a broader range of changes, offering deeper insights into the complexities of dynamic network analysis.
Our work is also related to the community recovery problem over censored block models. The semidefinite relaxation based method is a widely used approach to estimate the underlying community structures from adjacency matrices. The exact recovery (necessary and sufficient) conditions have been studied in [17, 18]. In our work, we assume that the community after the change is unknown, and thus a joint estimation and detection scheme is utilized, in which the semidefinite relaxation is used to estimate the community structure in a sequential manner. Furthermore, we derive new information-theoretic conditions for exact recovery under edge DP constraints.
Notation
Boldface uppercase letters denote matrices (e.g., ${\mathbf{A}}$), while boldface lowercase letters are used for vectors (e.g., ${\mathbf{a}}$). We denote scalars by non-boldface lowercase letters (e.g., $x$), and sets by capital calligraphic letters (e.g., ${\mathcal{X}}$). Let $[n]\triangleq[1,2,\cdots,n]$ represent the set of all integers from $1$ to $n$. We use $\operatorname{Bern}(p)$ to denote a Bernoulli random variable with success probability $p$. For asymptotic analysis, we say the function $f(n)=o(g(n))$ when $\lim_{n\rightarrow\infty}f(n)/g(n)=0$. Also, $f(n)=\mathcal{O}(g(n))$ means there exist some constant $C>0$ such that $|f(n)/g(n)|\leq C$, $\forall n$, and $f(n)=\Omega(g(n))$ means there exists some constant $c>0$ such that $|f(n)/g(n)|\geq c$, $\forall n$.
2 Preliminaries
In this section, we first introduce the binary censored block model used to model real-world network structures in Section 2.1. We present useful preliminaries and background on community estimation using maximum likelihood estimator and the corresponding semidefinite relaxation in Section 2.2. The definition of the notion of edge differential privacy used in this work is introduced in Section 2.3.
2.1 Binary Censored Block Model
We consider an undirected graph $G=({V},E)$ consisting of $n$ vertices, where the vertices are divided into two communities $C_{1}$ and $C_{2}$, and ${V}=C_{1}\cup C_{2}$. The community label for vertex $i$ is denoted by $\sigma^{*}_{i}\in\{-1,+1\}$, $\forall i\in[n]$ and we denote the label vector as $\bm{\sigma}^{\ast}=\{\sigma_{1}^{\ast},\ldots,\sigma_{n}^{\ast}\}$. Further, we assume that the graph $G$ is generated through a CBM. The connections between vertices are represented by the adjacency matrix $\mathbf{A}$, where each edge between two vertices $i<j$ is labeled as $A_{i,j}\in\{+1,-1,0\}$ and drawn as follows:
$A_{i,j}\!\sim\!\begin{cases}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0},&\text{if }\sigma^{*}_{i}=\sigma^{*}_{j},\\ p(1-\zeta)\delta_{-1}+p\zeta\delta_{+1}+(1-p)\delta_{0},&\text{if }\sigma^{*}_{i}\neq\sigma^{*}_{j},\end{cases}$ (1)
where $p$ represents the revealed probability of an edge. The parameter $\zeta$ lies within the interval $(0,\frac{1}{2})$, and $\delta$ denotes the Dirac delta function, which is centered around its subscript indicating the precise value of a parameter or a variable it is associated with. In our analysis, we consider an undirected graph without self-loops, i.e., $A_{i,i}=0$ for all $i$, ensuring no vertex is connected to itself, and $A_{i,j}=A_{j,i}$ for all $i\neq j$. For simplicity we denote this model as ${\mathsf{CBM}}(\bm{\sigma}^{\ast},p,\zeta)$. In this work, we focus on the dense regime of the CBM, parameterized by $p=a\log(n)/n$ for a positive constant $a$.
The goal of the community detection problem is to design a (stochastic) estimator $\hat{\bm{\sigma}}:\mathbf{A}\rightarrow\{-1,+1\}^{n}$ for community recovery (i.e., estimating the true labeling vector $\bm{\sigma}^{*}$) upon observing the adjacency matrix. We next define the notion of asymptotic exact recovery as a measure of performance of an estimator. Before we introduce the notion, we first present the discrepancy measure between any estimated community labeling vector $\hat{\bm{\sigma}}$ and the ground truth labels $\bm{\sigma}^{*}$ as
$\operatorname{err}(\hat{\bm{\sigma}},{\bm{\sigma}^{*}})$ $=\min\{\operatorname{Ham}(\hat{\bm{\sigma}},{\bm{\sigma}^{*}}),\operatorname{Ham}(-\hat{\bm{\sigma}},{\bm{\sigma}^{*}})\},$ (2)
where $\operatorname{Ham}(\cdot,\cdot)$ denotes the Hamming distance.
Definition 1 (Exact Recovery).
An estimator $\hat{\bm{\sigma}}(\mathbf{A})=\{\hat{\sigma}_{1},\hat{\sigma}_{2},\cdots,\hat{\sigma}_{n}\}$ satisfies exact recovery (up to a global flip) if the probability of error behaves as
$\operatorname{Pr}(\operatorname{err}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})=0)=1-o(1),$ (3)
where the probability is taken over both the randomness of the adjacency matrix ${\mathbf{A}}$ of the observed graph $G$ as well as the stochastic estimation process.
2.2 Maximum Likelihood Estimator and Semidefinite Relaxation
Under the ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ model, given an observation ${\mathbf{A}}$, the log-likelihood function can be calculated as (see Appendix A):
$\log p({\mathbf{A}}|\bm{\sigma})=$ $\left(\frac{1}{4}\log\frac{1-\zeta}{\zeta}\right)\bm{\sigma}^{T}{\mathbf{A}}\bm{\sigma}+{n\choose 2}\log(1-p)+E\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right),$ (4)
where $E=\sum_{i<j}|A_{i,j}|$ denotes the total number of edges for the adjacency matrix ${\mathbf{A}}$.
Therefore, for the binary CBM, given that $\zeta\in(0,\frac{1}{2})$, the Maximum Likelihood Estimator (MLE) of the label vector is the solution to the following optimization problem:
$\hat{\bm{\sigma}}_{\text{ML}}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}:\sigma_{i}=\pm 1\}.$ (5)
Information theoretically, it has been shown in [18] that for CBM within the dense regime, i.e., $p={a\log(n)}/{n}$, exact recovery is possible if and only if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$.
It is worth mentioning that the MLE is equivalent to solving the max-cut problem in the graph, which is known to be NP-hard [18]. Subsequently, several works in the literature have proposed polynomial time algorithms that can achieve the exact recovery. For instance, it has been shown in [18] that semidefinite relaxation of the MLE can also achieve the optimal recovery threshold. We next present the optimization problem of the MLE via semidefinite relaxation as proposed in [18].
Let us first define $\mathbf{Y}={\bm{\sigma}}{\bm{\sigma}}^{T}$, where $Y_{i,i}=1,\forall i\in[n]$. The semidefinite relaxation approach solves the following optimization problem:
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\hskip 7.22743pt\text{tr}({\mathbf{A}}\mathbf{Y})$
s.t. $\hskip 7.22743pt\mathbf{Y}\succcurlyeq\mathbf{0},\,Y_{i,i}=1,\forall i\in[n].$ (6)
It has been shown in [18] that the MLE obtained via semidefinite programming achieves exact recovery (after rounding the solution of (6)) if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$.
2.3 Differential Privacy
In addition to community recovery, we require that the recovery algorithm for community detection also protects the individual connections (i.e., the edges in the graph $G$) in the network.
Privacy Threat Model: In the context of community detection within graph data, our threat model is structured around the principle of edge differential privacy (DP), incorporating both local and central settings. Local DP offers robust privacy guarantees during the data collection process, ensuring that individual data contributions are privatized at the source. This is crucial in scenarios where trust in the data curator is limited. Conversely, in the central setting, privacy protections are applied to the aggregated data, focusing on preserving the presence or absence of single edges. The primary sensitivity in this model lies in the relationships between nodes rather than the nodes’ labels. This sensitivity arises from potential privacy risks of exposing how individuals or entities are interconnected, which could reveal undisclosed or sensitive information. The model assumes that an adversary could have partial knowledge about the graph’s structure and seeks to infer additional undisclosed relationships. Our approach aims to ensure that the detection of communities within this graph does not compromise the privacy of the relationships between the nodes in both local and central settings, allowing for meaningful analysis of the overall graph structure while safeguarding individual privacy.
Similar to [22], we consider the privacy mechanisms where each privatized edge $\tilde{\mathbf{A}}_{i,j}$ will depend on the raw edge ${\mathbf{A}}_{i,j}$ through a conditional distribution $Q$, i.e., $\tilde{\mathbf{A}}_{i,j}|{\mathbf{A}}_{i,j}=a\sim Q(\cdot|a)$.
Definition 2 ($\epsilon$-edge LDP).
We say the privacy mechanism $\tilde{A}_{i,j}\sim Q(\cdot|A_{i,j})$, with $\tilde{\mathbf{A}}$ being the privatized and ${\mathbf{A}}$ the original adjacency matrix, is $\epsilon$-edge LDP, if for all $i<j$ and all $a,a^{\prime},\tilde{a}\in\{-1,+1,0\}$, it holds that
$\frac{Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a)}{Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a^{\prime})}\leq e^{\epsilon}.$
We next present the notion of $\epsilon$-edge central differential privacy (CDP) [20].
Definition 3 ($\epsilon$-edge CDP).
An estimator $\hat{\bm{\sigma}}$ satisfies $\epsilon$-edge CDP for some ϵ+italic-$\epsilon\in\mathds{R}^{+}$, if for all pairs of adjacency matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$ that differ in one edge, we have
$\frac{\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})=\bm{\sigma})}{\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A}^{\prime})=\bm{\sigma})}\leq e^{\epsilon},\,\forall\bm{\sigma},$ (7)
where the probabilities are computed only over the randomness in the estimation process.
It is worthwhile mentioning that Central DP is a weaker guarantee compared to the local DP, as it can be readily shown that the local DP guarantee implies central DP.
3 Online Community Change Detection
3.1 Problem Setup
Suppose we observe a sequence of undirected graphs $\{G_{t},t\in\mathbb{N}\}$ generated through the ${\mathsf{CBM}}(\bm{\sigma}_{t}^{*},p,\zeta)$ model, with $\mathbf{A}_{t}$ being the adjacency matrix of $G_{t}$, as detailed in Section 2.1. We consider the setting with known $p,\zeta$ and dynamic communities in which the true labeling vector $\bm{\sigma}_{t}^{*}$ can be time-varying. This is commonly seen in practical applications. For example, in social network data, the users may switch communities due to certain events [38].
We consider a special case, in which there is an abrupt change in the community structure at time $\nu$, which is called a change-point and takes a deterministic but unknown value. The problem can be formulated as the following hypothesis test:
$\begin{array}[]{lll}H_{0}:&\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{pre}},&~{}~{}~{}t=1,2,\dots\\ H_{1}:&\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{pre}},&~{}~{}~{}t=1,\dots,\nu-1,\\ &\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{post}},&~{}~{}~{}t=\nu,\tau+1,\dots\\ \end{array}$ (8)
where $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ represent the pre- and post-change community labels, respectively. We further assume $\bm{\sigma}^{\text{pre}}$ is known since it can usually be estimated from historical data collected under the pre-change (nominal) regime. Let $\mathbb{P}_{\nu}$ denote the probability measure on the observation sequence when the change-point is $\nu$, and let $\mathbb{E}_{\nu}$ denote the corresponding expectation. Also, $\mathbb{P}_{\infty}$ and $\mathbb{E}_{\infty}$ denote the probability and expectation operator when there is no change (i.e., $\nu=\infty$). For brevity, we write $\mathbb{P}$ and $\mathbb{E}$ as the probability and expectation operator when $\nu=1$.
We make the following assumptions on the change in community structures.
Assumption 1.
There exists a fixed integer $\Delta\in(0,n/2)$ such that the Hamming distance between the pre-change community label $\bm{\sigma}^{\text{pre}}$ and post-change label $\bm{\sigma}^{\text{post}}$ satisfies
$\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})\leq\Delta.$ (9)
When the total number of nodes $n$ is large and $\Delta$ is much smaller than $n/2$, Assumption 1 implies that the change in the community structure is sparse.
3.2 Performance Criteria
Our goal is to design a detection algorithm that is able to protect the user privacy as in Definition 2 while detecting the change as quickly as possible from streaming data. The detection is performed through a stopping time $\tau$ on the observation sequence at which the change is declared [29].
More specifically, we introduce two performance measures commonly used in online detection problems: detection delay and false alarm performances. The commonly adopted definition of Worst-case Average Detection Delay (WADD) is [23]
${\mathsf{WADD}}(\tau)=\underset{\nu\geq 1}{\sup}\operatorname*{ess\,sup}\ \mathbb{E}_{\nu}\left[(\tau-\nu+1)^{+}|\mathcal{F}_{\nu-1}\right],$ (10)
where $\{\mathcal{F}_{k},\ k\in\mathbb{N}\}$ is the filtration $\mathcal{F}_{k}=\sigma(G_{1},\ldots,G_{k})$, with $\mathcal{F}_{0}$ denoting the trivial sigma algebra.
We measure the false alarm performance of a detector (stopping time) $\tau$ in terms of its mean time to false alarm $\mathbb{E}_{\infty}\left[\tau\right]$, defined as Average Run Length (ARL). And we denote by $\mathcal{C}(\gamma)$ the set of all tests for which the ARL is at least $\gamma$, i.e.,
$\mathcal{C}(\gamma)=\left\{\tau:\;\mathbb{E}_{\infty}\left[\tau\right]\geq\gamma\right\}.$ (11)
The optimization problem of interest is
$\inf_{\tau\in\mathcal{C}(\gamma)}{\mathsf{WADD}}(\tau).$ (12)
Lorden showed that Page’s Cumulative Sum (CUSUM) test [28] solves the problem in (12) asymptotically; this result was strengthened to exact optimality in [26]. Under the private setting, we aim to design a detection method that attains both the data privatization and efficient change detection.
4 Proposed Method: Private Window-Limited CUSUM Test
In this section, we introduce the proposed joint estimation and detection scheme, in which we perform sequential estimates for the unknown post-change community structure $\bm{\sigma}^{\text{post}}$ and use it to calculate the detection statistic. We first introduce the private mechanism that outputs privatized data through a conditional distribution that is conditional on raw data. The proposed method is summarized in Algorithm 1.
Algorithm 1 Private Window-Limited CUSUM Test for Online Community Change Detection
  Input: Sequence of adjacency matrices $\{{\mathbf{A}}_{t},t\in\mathbb{N}\}$, window size $w$, threshold $b$.
  Private mechanism: for each $A_{t,i,j}$, output the privatized sample
$\operatorname{Pr}(\tilde{A}_{t,i,j}=A_{t,i,j})$ $=\frac{e^{\epsilon}}{e^{\epsilon}+2},$
$\operatorname{Pr}(\tilde{A}_{t,i,j}=x)$ $=\frac{1}{e^{\epsilon}+2},\quad\forall x\neq A_{t,i,j}.$
  Initialize $S_{1}=\cdots=S_{w}=0$; $t=w+1$.
  repeat
     Calculate $\bm{\hat{\sigma}}_{t-1}$ via solving (16).
     Update $S_{t}=(S_{t-1})^{+}+\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})$.
     if $S_{t}\geq b$ then
        Raise Alarm
     end if
     $t=t+1$;
  until Raise Alarm
4.1 Privacy Mechanism
In this subsection, we present the graph Perturbation Mechanism that we adopted for private community detection. Furthermore, we present and analyze a noise-graph addition mechanism based on Warner’s randomized response [35]. The basic idea is to perturb the edges of the random graph (i.e., the adjacency matrix $\mathbf{A}$) to obtain a perturbed adjacency matrix $\tilde{\mathbf{A}}$ of $\mathbf{A}$, where each element $A_{i,j}$ is perturbed independently to satisfy $\epsilon$-edge LDP.
More specifically, when $A_{i,j}=x$ where $x\in\{0,+1,-1\}$, the mechanism outputs $\tilde{A}_{i,j}\sim Q_{x}(\cdot)$. We adopt a special case for graph perturbation mechanism where $Q_{x}(y)=c_{2}=\frac{1}{e^{\epsilon}+2}$ when $x\neq y$, and $Q_{x}(y)=c_{1}=1-2c_{2}=\frac{e^{\epsilon}}{e^{\epsilon}+2}$ when $x=y$. Note that for the original data distribution ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, the perturbed elements in the adjacency matrix $\tilde{\mathbf{A}}$ are distributed as ${\mathsf{CBM}}(\bm{\sigma},\tilde{p},\tilde{\zeta})$ with (see Appendix E),
$\tilde{p}=\frac{2+p(e^{\epsilon}-1)}{e^{\epsilon}+2},\ \tilde{\zeta}=\frac{1+p\zeta(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}.$ (13)
Also, note that the randomization mechanism is structured in such a way that $\tilde{A}_{i,j}=\tilde{A}_{j,i}$ and $\tilde{A}_{i,i}=0$.
4.2 Detection Procedure
We apply the detection procedure on the privatized data $\{\tilde{\mathbf{A}}_{t},t\in\mathbb{N}\}$. We adopt the recently proposed window-limited CUSUM test [36], with modifications tailored for the community estimation problem. At each time $t$, we use the current community structure estimate $\bm{\hat{\sigma}}_{t-1}$ obtained from most recent private samples $\{\tilde{\mathbf{A}}_{t-w},\ldots,\tilde{\mathbf{A}}_{t-1}\}$. Then we calculate the detection statistics recursively via
$S_{t}=(S_{t-1})^{+}+\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1}),\,t\geq 1,$ (14)
with initial value $S_{0}=0$, and $(x)^{+}:=\max\{x,0\}$. In (14), the function $\ell$ is defined as
$\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma},\bm{\sigma}^{\prime}):=\log\frac{p(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\prime})}{p(\tilde{\mathbf{A}}_{t};\bm{\sigma})},$ (15)
which calculates the log-likelihood ratio for each randomized sample $\tilde{\mathbf{A}}_{t}$ and is an essential element in the detection algorithms. It is worthwhile mentioning that, by definition, the estimate $\bm{\hat{\sigma}}_{t-1}$ is independent with the current sample $\tilde{\mathbf{A}}_{t}$. Such independence is essential for providing the theoretical guarantee for the detection procedure.
We then introduce how we obtain the estimator $\bm{\hat{\sigma}}_{t-1}$. We adopt the maximum likelihood estimator for windowed samples:
$\bm{\hat{\sigma}}_{t}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\left(\sum_{i=0}^{w-1}\tilde{\mathbf{A}}_{t-i}\right)\bm{\sigma}\}.$
Following Section 2.2, we define $\mathbf{Y}={\bm{\sigma}}{\bm{\sigma}}^{T}$ and use the Semidefinite Relaxation for solving $\bm{\sigma}$ (after rounding the solution of (16)):
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\ \text{tr}(\left(\sum_{i=0}^{w-1}\tilde{\mathbf{A}}_{t-i}\right)\mathbf{Y})$ (16)
s.t. $\ \mathbf{Y}\succcurlyeq\mathbf{0},Y_{i,i}=1,\forall i\in[n].$
We note that the above optimization problem can be viewed as the approximate maximum likelihood solution to a multi-view stochastic block model [39]. Moreover, it is worthwhile mentioning that we allow great flexibility in the choice of window size $w$ here. When choosing $w=1$, we use the randomized sample $\tilde{\mathbf{A}}_{t-1}$ to estimate the current community structure and then accumulate the estimated log-likelihood ratio on sample $\tilde{\mathbf{A}}_{t}$. We leave more involved discussion on the impact of window size for future work.
Given the detection statistic $S_{t}$, the detection procedure is conducted via a stopping rule,
$T:=\inf\{t:S_{t}\geq b\},$ (17)
where $b$ is a pre-specified detection threshold chosen to satisfy the desired ARL constraint (i.e., such that $T\in\mathcal{C}(\gamma)$). We will discuss the choice of the detection threshold $b$ in Section 5.
5 Theoretical Analysis
In this section, we present the theoretical analysis for the proposed private detection procedure. We first show that the procedure is indeed $\epsilon$-edge LDP (Lemma 5.1). Then, we present theoretical guarantees for the average run length, and the worst-case average detection delay. We also compare with the information-theoretic lower bound on the detection delay in the non-private scenario (Theorem 5.1) to demonstrate the tradeoff between privacy constraints and the detection performance.
By the construction of the private mechanism, it is obvious we have the following guarantee on LDP.
Lemma 5.1 (Privacy guarantee).
The perturbation used in Section 4.1 satisfies $\epsilon$-edge LDP.
Remark 5.1.
Lemma 5.1 offers a privacy guarantee for each individual release. To evaluate the overall privacy leakage when the detection algorithm is applied to $w$ adjacency matrices, we utilize the composition property of DP. Notably, advanced composition techniques based on Rényi divergence can be employed to optimize the total privacy leakage. Hence, our analysis concludes that the cumulative privacy leakage for releasing $w$ adjacency matrices per user is bounded by $\mathcal{O}(\sqrt{w}\epsilon)$ in the context of edge DP. This paper, however, is primarily concerned with exploring the privacy ramifications for each single adjacency matrix release, rather than the cumulative effect over multiple releases.
We first present the well-known information-theoretic lower bound on the detection delay for online change detection on the above privatized data. Such lower bound will serve as the baseline, and we will compare the performance of our proposed method with such lower bound both theoretically and numerically. We denote the KL divergence between the post- and the pre-change privatized distributions as
$\tilde{I}_{0}:=\tilde{I}(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}}):=\mathbb{E}[\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})],$ (18)
where ${\mathbb{E}}$ is the expectation under the post-change regime as defined above, and $\tilde{\mathbf{A}}$ is the privatized data. The known information-theoretic lower bound on the detection delay on the privatized data is a function of the KL divergence as shown below.
Theorem 5.1 (Information-Theoretic Lower Bound on WADD [32]).
A lower bound on the WADD of any test, applied on the privatized sequence of adjacency matrices $\{\tilde{\mathbf{A}}_{t},t\in\mathbb{N}\}$, that satisfies the average run length constraint $\mathbb{E}_{\infty}\left[\tau\right]\geq\gamma$ is
${\mathsf{WADD}}(\tau)\geq\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1)).$
Remark 5.2.
Here we note that for given distribution parameter $p,\zeta$, the KL divergence $I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})$ between the post-change ${\mathsf{CBM}}(\bm{\sigma}^{\text{post}},p,\zeta)$ and pre-change ${\mathsf{CBM}}(\bm{\sigma}^{\text{pre}},p,\zeta)$ equals to
$\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\right].$
where $C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}=\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}$. See Appendix A for detailed derivations.
5.1 Performance Analysis for the Proposed Method
In the proposed detection method, we have a sliding window estimate for the unknown community structures. We first analyze the sufficient conditions for exact community recovery under privacy constraints in Section 5.1.1.
5.1.1 Sufficient Condition for Exact Community Recovery
Theorem 5.2.
The graph perturbation based recovery algorithm via Warner’s randomized response satisfies $\epsilon$-edge CDP, $\forall\epsilon\geq\epsilon_{n}=\Omega(\log(n))$, $\forall a>\frac{2(n^{3/2}-n)}{(n-1)\log(n)}$, and it satisfies exact recovery if
$a\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}>\left(\frac{\sqrt{n}}{\sqrt{n}-1}\right)\times\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right).$ (19)
The proof is given in Appendix E.
Remark 5.3.
Observe that, if $\epsilon$ remains constant while $n$ increases, it follows that from (13), $\lim_{n\rightarrow\infty}\tilde{\zeta}=1/2$. This suggests that with a consistent $\epsilon$, the characteristics of both inter- and intra-community edges converge in the asymptotic limit, rendering exact recovery infeasible. However, Theorem 5.2 demonstrates that exact recovery can be achieved when the leakage grows logarithmically with $n$.
5.1.2 Necessary Condition for Exact Community Recovery
We subsequently introduce an information-theoretic lower bound that delineates the minimum necessary privacy levels for achieving exact community recovery in $\epsilon$-edge CDP frameworks under CBMs. This lower bound serves as a critical benchmark for assessing the effectiveness and limitations of privacy-preserving mechanisms in precise community detection within graph structures, particularly under the stringent conditions imposed by $\epsilon$-edge CDP.
Theorem 5.3 (Converse Result).
Suppose there exists an $\epsilon$-edge CDP mechanism such that for any labeling vector $\bm{\sigma}\in\{\pm 1\}^{n}$, on $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, output $\hat{\bm{\sigma}}\in\{\pm 1\}^{n}$ satisfying (3). Then, the privacy leakage must scale at least as
$\epsilon$ $\geq\frac{1}{2}\log\left[1+\frac{2\log(n)-\log(8e)}{p^{\prime}(4n-32)}\right],$ (20)
where $p^{\prime}=2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1$, $p=a\log(n)/n$ and $\zeta\in(0,1/2)$.
The proof is given in Appendix F.
Remark 5.4.
The lower bound derived in this context reveals the direct connection between the CBM parameters $a$ and $\zeta$, and the privacy budget $\epsilon$ required for exact community recovery. Distinctly, this lower bound decays slower than the bound derived from the traditional packing lower bound argument method, as discussed in [33]. In contrast to the constant nature of our derived lower bound, the packing argument suggests a more dynamic relationship, specifically, $\epsilon=\Omega\left({\log(n)}/{{n\choose 2}}\right)$, highlighting a less robust link with $n$. For smaller values of $\epsilon\leq 1$, the lower bound on the privacy budget in (20) scales as $\epsilon=\Omega\left({\log(n)}/{n}\right)$, which indicates a nuanced scaling behavior in comparison to the previously discussed bound. It is worth highlighting that the proposed graph perturbation mechanism requires $\epsilon=\Omega(\log(n))$. There remains an interesting opportunity to close this gap by devising more sophisticated mechanisms.
These results provide us sufficient and necessary conditions on the privacy level $\epsilon$ in order to guarantee efficient detection using privatized data.
5.1.3 False Alarm and Detection Delay
Under the exact recovery, we analyze the property of the estimated $\hat{\bm{\sigma}}$ and then the detection performance. All technical details can be found in Appendix G.
First of all, we have the following guarantee for the average run length that measures the false alarm performance.
Lemma 5.2 (Average Run Length).
The proposed detection method (17) satisfies, for any detection threshold $b$,
${\mathbb{E}}_{\infty}[T]\geq e^{b}.$
It is obvious to see that we can choose the detection threshold $b=\log\gamma$ such that the average run length constraint ${\mathbb{E}}_{\infty}[T]\geq\gamma$ is satisfied. Therefore, we adopt this threshold in subsequent delay analysis, and provide an upper bound for the detection delay in the following.
Theorem 5.4 (WADD Upper Bound).
Assume the exact recovery condition (19) is satisfied, we have the following upper bound on the worst-case performance of the detection procedure (17), when choosing $b=\log\gamma$, as $\gamma\to\infty$, and $w\tilde{I}_{0}=o(\log\gamma)$:
${\mathbb{E}}_{0}[T]\leq\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1)),$ (21)
where $\tilde{I}_{0}$ is the KL divergence of the privatized post-change distribution from the privatized pre-change distribution defined in (18).
Remark 5.5.
We note that, compared with the results in [36], our setting and method here have two fundamental differences. First, the parameter to be estimated in our case, $\bm{\sigma}$, is discrete, while the maximum likelihood estimate in [36] was conducted for continuous values. Second, we use an approximate maximum likelihood estimator resulting from the semidefinite relaxations, whereas [36] employed the exact maximum likelihood estimator. These differences call for new analysis techniques to address discrete parameter spaces and the approximation error of semidefinite relaxations, which we provide in the appendix.
5.2 Delay Lower Bound with Private Constraint
Our objective is to detect changes in dynamic networks under privacy constraints. The lower bound on the detection delay in Theorem 5.1 is the information-theoretic lower bound under the particular private mechanism used in Section 4.1. In this section, we further investigate the minimax lower bound for arbitrary private algorithms that satisfies $\epsilon$-edge LDP, as defined as
$\inf_{Q\in\mathcal{Q}_{\epsilon}}\inf_{\tau\in\mathcal{C}(\gamma)}\underset{\nu\geq 1}{\sup}\operatorname*{ess\,sup}\ \mathbb{E}_{\nu,Q}\left[(\tau-\nu+1)^{+}|\mathcal{F}_{\nu-1}\right],$ (22)
where $\mathcal{Q}_{\epsilon}$ is the set of all privacy mechanisms that satisfy the $\epsilon$-edge LDP constraint, $\tau$ ranges over all stopping rules in the set $\mathcal{C}(\gamma)$ satisfying the ARL constraint; and the inner supremum denotes the worst-case detection delay over all possible change-point locations, with $\mathbb{E}_{\cdot,Q}$ denoting the expectation under the current privacy mechanism $Q$. We aim to provide a general lower bound to (22) and compare with the delay of our proposed method in the end.
To solve the above minimax lower bound, we first analyze the KL divergence after the privacy mechanism. We denote $p_{\infty}({\mathbf{A}})$ and $p_{0}({\mathbf{A}})$ as the probability distribution for the adjacency matrix ${\mathbf{A}}$ in the pre-change and post-change regime, respectively. And denote $\tilde{p}_{\infty}({\mathbf{A}})$ and $\tilde{p}_{0}({\mathbf{A}})$ as the corresponding two distributions for the privatized data distribution. Note that we have
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})$ $=\mathbb{E}_{\tilde{p}_{0}}\bigg{[}\sum_{1\leq i<j\leq n}\log\frac{\tilde{p}_{0}(\tilde{A}_{ij})}{\tilde{p}_{\infty}(\tilde{A}_{ij})}\bigg{]}.$
Denote $q_{1}(z)=q(\tilde{A}_{ij}=z|A_{ij}=1)$ as the probability that the privatized edge equals to $z$ when the raw edge is $+1$, and similarly we define $q_{0}(z)$ and $q_{-1}(z)$. By the privacy constraint we have $\frac{q_{x}(z)}{q_{y}(z)}\in[e^{-\epsilon},e^{\epsilon}]$ for any $x,y\in\{-1,0,+1\}$. Based on this, and following [11], we have the following upper bound for the KL divergence after any $\epsilon$-edge LDP perturbation. Technical details can be found in Appendix H.
Lemma 5.3 (KL divergence after perturbation).
For any privacy mechanism $Q$ satisfying the $\epsilon$-edge LDP in Definition 2, we have the KL divergence after perturbation is upper bounded as follows:
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})\leq\min\{4,e^{2\epsilon}\}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}\bigg{[}{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\bigg{]}.$
The above Lemma implies that
$\inf_{Q\in\mathcal{Q}_{\epsilon}}\inf_{\tau\in\mathcal{C}(\gamma)}\underset{\nu\geq 1}{\sup}\operatorname*{ess\,sup}\ \mathbb{E}_{\nu,Q}\left[(\tau-\nu+1)^{+}|\mathcal{F}_{\nu-1}\right]\geq\\ \frac{\log\gamma}{\min\{4,e^{2\epsilon}\}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})}(1+o(1)).$ (23)
Meanwhile, recall that in our privacy mechanism, we have that the KL divergence after the randomized perturbation becomes $\tilde{I}_{0}=\left(\log\frac{1+p(1-\zeta)(e^{\epsilon}-1)}{1+p\zeta(e^{\epsilon}-1)}\right)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})\approx p(1-2\zeta)(e^{\epsilon}-1)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})$ for small $\epsilon$ (i.e., strong privacy constraint), which matches with the denominator in (23) up to a constant factor.
6 Numerical Results
In this section, we present numerical results to assess the performance of our proposed algorithm. The proposed mechanisms are implemented in MATLAB, and the optimization (via SDP method) is done through the CVX solver [15].
6.1 Simulation Study
In the numerical results, we evaluate two performance metrics: 1) the detection delay as defined in (10); and 2) the average run length defined as $\mathbb{E}_{\infty}[\tau]$. In our numerical experiments, we simulate the average detection delay as $\mathbb{E}_{0}[\tau]$ by setting the change-point at time 00.
We set the number of nodes $n=50$, and select the pre- and post-change community label vector $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ such that $\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})=2$. The distribution parameter is $p=a\log(n)/n$, and $\zeta=0.1$. The private parameter is set as $\epsilon$. We consider two senarios: (1) Case 1 ($a=5,\epsilon=1.5$) represents the case that condition (19) is satisfied, and we use the smallest window size $w=1$ in such scenario; (2) Case 2 ($a=6,\epsilon=1$) represents the case that condition (19) is violated and the community recovery can be unstable, thus we use multi-window with $w=10$.
We demonstrate the sample trajectories of the detection statistics in Figure 1; we observe that the detection statistics under Case 1 are more stable under the pre-change regime (mostly below zero) and become mostly positive after the change, yielding a relatively small delay as visualized in Figure 2 (Left). We also compare the simulated delay of our proposed method with the lower-bound in Theorem 5.1 in Figure 2 (Left); it can be seen that the delay of our method is close to the lower-bound and matches the lower-bound in the first order. It is worth mentioning that the KL divergence between the post- and pre-change distributions in Case 1 is significant. Thus, the information-theoretic lower bound on the delay is very small, and our proposed method can also obtain a delay of less than four, even for large ARL constraints.
Moreover, we visualize the trajectory for Case 2 in Figure 1 (Right), and the pre-change statistics become more noisy due to the violation of condition (19), while we are still able to detect the change, with a relatively larger delay as shown in Figure 2 (Right), which of course is no longer near-optimal. We want to comment that this shows even if the condition (19) is violated, the proposed detection statistics are still able to detect the change, and we left the more involved theoretical analysis on the choice of window size $w$ and to further improve the detection performance as future work.
Refer to caption Refer to caption
Case 1 Case 2
Figure 1: Sample trajectories of the detection statistics.
Refer to caption Refer to caption
Case 1 Case 2
Figure 2: Detection Delay v.s Average Run Length.
6.2 Real Case Study
We implemented the online detection algorithm on the worldwide agricultural trade network dataset, which is publicly available at the Food and Agricultural Organization of the United Nations (The dataset is available at https://www.fao.org/faostat/en/#data/TM). We select annual data from year 1991 to 2015. The dataset includes the quantity and value of all food and agricultural products imported/exported annually by all the countries in the world, and it has been previously analyzed in studies such as [34].
We select $n=50$ countries with the most significant importing/exporting relationship to construct the adjacency matrices. For each year, the entry $A_{i,j}$ in the network is set to $A_{i,j}=0$ if there is no import-export relation between two countries, $A_{i,j}=1$ if the total trade column of the top $4$ agricultural products exceeds a pre-specified threshold, and $A_{i,j}=-1$ otherwise. Consequently, each node corresponds to a country, and each edge represents the export-import relationship between countries. We use the data of the year 1991 to estimate the pre-change community label, parameters $p$, and $\zeta$.
Refer to caption Refer to caption Refer to caption Refer to caption
Raw Network $\epsilon=1.5$ $\epsilon=1$ $\epsilon=0.8$
Figure 3: Results on the agricultural trade dataset.
Figure 3 is the plot of the resulting CUSUM detection statistics, which matches well with the change-point 2002 identified by previous studies, such as [34]. As mentioned in [34], the year 2002 was a significant low and turning point in global commodity prices, which likely influenced trade patterns since as countries adjusted their import and export practices in response to changing prices. From the trajectory of detection statistics after applying randomized perturbation with different levels of $\epsilon$ values, we can see the tradeoff between privacy and utility (detection effectiveness): stronger privacy constraints (lower $\epsilon$ values) make it more challenging to detect the change-point.
Furthermore, we also compare the community estimate results using the data in the years 1991 (before the change) and 2003 (after the detected change-point). The estimated community labels for seven countries (United Arab Emirates, Egypt, Hungary, Lebanon, Viet Nam, Saudi Arabia, Romania) changed from $-1$ to $1$. Here, $1$ represents the larger community at both times, indicating that countries such as “Viet Nam” have become more engaged with other leading countries in the global market for agricultural imports and exports after 2002.
7 Conclusion and Discussion
Building on these findings, we identify several avenues for future research. Firstly, we may consider the detection of general changes, including changes in both the community label vector $\bm{\sigma}$ and the parameters $p,\zeta$ of the CBM model. Changes in $p$ and $\zeta$ will represent practical scenarios in which the local connectivity within a community is strengthened or weakened after the change. Secondly, we leave the more involved discussion on the impact of the multi-view stochastic model [39] on our joint detection and estimation scheme, as well as the corresponding impact of the window size $w$, for future investigation. We conjecture that there is an optimal tradeoff between the detection performance and computational/memory efficiency by carefully choosing the window size $w$, especially in the non-asymptotic regime where the network size $n$ is limited. Finally, this work opens numerous opportunities for private community detection problems, and more general differential privacy mechanisms such as the theoretical guarantee under different privacy can be studied for future work.
Appendix A Derivation of the Log-Likelihood and the KL Divergence
For a given graph $G_{t}$ at time $t$ with adjacency matrix ${\mathbf{A}}_{t}$, denote $E_{t,1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=1\}$ as the number of edges within the same community with edge label $1$, or in different communities with edge label $-1$; similarly $E_{t,-1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=-1\}$. Let $E_{t}=E_{t,1}+E_{t,-1}$ be the total number of edges at time $t$. Then we have the likelihood function is
$p(\mathbf{A};\bm{\sigma},p,\zeta)=[p(1-\zeta)]^{E_{t,1}}[p\zeta]^{E_{t,-1}}(1-p)^{{n\choose 2}-E_{t}}.$
Therefore the log-likelihood is
$\log p(\mathbf{A};\bm{\sigma},p,\zeta)=E_{t}\log\frac{p}{1-p}+{n\choose 2}\log(1-p)+E_{t,1}\log\frac{1-\zeta}{\zeta}+E_{t}\log\zeta.$
Note that
$\frac{1}{2}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}=E_{t,1}-E_{t,-1}\Rightarrow 2E_{t,1}-E_{t}=\frac{1}{2}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}\Rightarrow E_{t,1}=\frac{1}{4}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}+\frac{1}{2}E_{t}.$
Substitute this into the above log-likelihood we obtain
$\log p(\mathbf{A}_{t};\bm{\sigma},p,\zeta)=\left(\frac{1}{4}\log\frac{1-\zeta}{\zeta}\right)\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}+{n\choose 2}\log(1-p)+E_{t}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right).$ (24)
By the log-likelihood function we have that the MLE for $\bm{\sigma}$ is $\hat{\bm{\sigma}}_{\text{MLE}}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}\}$.
KL Divergence (Remark 5.2)
Based on (24), the KL divergence between ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and ${\mathsf{CBM}}(\bm{\sigma}^{\prime},p,\zeta)$ can be calculated as
$I$ $={\mathbb{E}}_{\bm{\sigma}^{\prime}}[\log p({\mathbf{A}};\bm{\sigma}^{\prime},p,\zeta)-\log p({\mathbf{A}};\bm{\sigma},p,\zeta)]$ (25)
$=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta){n\choose 2}+{n\choose 2}\log(1-p)+p{n\choose 2}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right)$
$\quad\quad-\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma}_{i}^{\prime}\bm{\sigma}_{j}^{\prime}-{n\choose 2}\log(1-p)-p{n\choose 2}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right)$
$=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma},\bm{\sigma}^{\prime}}\right],$
where $C_{\bm{\sigma},\bm{\sigma}^{\prime}}:=\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma}_{i}^{\prime}\bm{\sigma}_{j}^{\prime}$ represent the similarity level between the community label vectors $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ and it satisfies $|C_{\bm{\sigma},\bm{\sigma}^{\prime}}|\leq{n\choose 2}$.
Appendix B Ancillary Lemmas
We provide auxiliary results and definitions that are used throughout the proofs.
Lemma B.1 ([17]).
Let $\{X_{i}\}_{i=1}^{m}$ be a collection of independent random variables, where $p=a\frac{\log n}{n}$, $m=n+o(n)$ and each $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Then,
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg{)}\leq 1-n^{-a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}+o(1)},$ (26)
where $k_{n}=(1+o(1))\frac{\log(n)}{\log(\log(n))}$.
Lemma B.2.
(The Chernoff-Hoeffding Inequality [24]) For a binomial random variable $X$ with $n$ trials and mean $\mu_{X}$, the probability that $X$ deviates from the mean by more than $\beta n$ can be bounded as,
$\operatorname{Pr}(|X-\mu_{X}|\geq\beta n)\leq 2e^{-2\beta^{2}n}\triangleq\delta^{\prime},$ (27)
for any $\beta>0$, and any $\delta^{\prime}\in[0,1)$.
Lemma B.3.
(Bernstein’s Inequality [24]) Let $\{X_{i}\}_{i=1}^{n}$ be a collection of zero-mean independent random variables, where each $X_{i}$ is bounded by $M$, almost surely. Then, for any $t\geq 0$, we have
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{n}X_{i}>t\bigg{)}\leq\exp\left[-\frac{t^{2}}{\sum_{i=1}^{n}E(X_{i}^{2})+\frac{Mt}{3}}\right].$ (28)
We present a high probability result for upper bounding the spectral norm of $\mathbf{A}-\mathds{E}(\mathbf{A})$.
Lemma B.4.
(Matrix concentration inequality for dense regime [13]) For a matrix $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and $p=\Omega(\log(n)/n)$, we have
$\|\mathbf{A}-\mathds{E}(\mathbf{A})\|^{2}\leq c^{\prime}\log(n)$ (29)
w.h.p. for some constant $c^{\prime}$.
Appendix C Brief Refresher on Differential Privacy
Definition 4 ($(\varepsilon,\delta)$-DP).
A randomized algorithm $\mathcal{A}:\mathcal{D}\rightarrow\mathbb{R}^{d}$ satisfies $(\varepsilon,\delta)$-DP if for any two neighboring datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ in one record, and any measurable subset $\mathcal{S}\subseteq\text{Range}(\mathcal{A})$, we have
$\operatorname{Pr}(\mathcal{A}(\mathcal{D})\in\mathcal{S})\leq e^{\varepsilon}\operatorname{Pr}(\mathcal{A}(\mathcal{D}^{\prime})\in\mathcal{S})+\delta,$ (30)
where $\epsilon$ is coined as privacy budget and $\delta$ represents failure probability. It worth noting that smaller $\epsilon$, $\delta$ correspond to stronger privacy guarantee and vice versa. The setting when $\delta=0$ is referred as pure $\varepsilon$-DP.
Definition 5.
(Sequential Composition of DP [12]) If a randomized mechanism $\mathcal{M}_{i}$ satisfies $(\epsilon_{i},\delta_{i})$-DP for $i=1,2,\cdots,n$, then for any $\delta_{i}\in(0,1]$, their composition $\mathcal{M}$ defined by $\mathcal{M}\triangleq(\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{T})$ satisfies $(\sum_{i=1}^{T}\epsilon_{i},\sum_{i=1}^{T}\delta_{i})$-DP.
Definition 6.
(Advanced Composition of DP [12]) If a sequence of randomized mechanisms $\mathcal{M}_{i}$ each satisfies $(\epsilon,\delta)$-differential privacy for $i=1,2,\cdots,T$, then their composition $\mathcal{M}$, defined by $\mathcal{M}\triangleq(\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{T})$, satisfies $\left(\sqrt{2T\ln(1/\delta^{\prime})}\epsilon+T\epsilon(e^{\epsilon}-1),T\delta+\delta^{\prime}\right)$-differential privacy for any $\delta^{\prime}>0$.
Appendix D Deriving Recovery Condition for Exact Recovery
In this subsection, we analyze the error probability of the semidefinite relaxation of the ML estimator as shown in eqn. (6). Moreover, we derive the necessary conditions for the optimal solution. The Lagrangian function is written as follows:
$\mathcal{L}({\mathbf{A}},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}({\mathbf{A}}\mathbf{Y})+\operatorname{tr}(\mathbf{S}\mathbf{Y})-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$ (31)
where $\mathbf{S}\succcurlyeq\mathbf{0}$ and $\mathbf{D}=\operatorname{diag}(d_{i})$. Then,
$\nabla_{\mathbf{Y}}\mathcal{L}={\mathbf{A}}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$ (32)
In order to satisfy the first order stationary condition, we must satisfy the following (deterministic) condition:
$\mathbf{S}^{*}=\mathbf{D}^{*}-{\mathbf{A}},$ (33)
From the KKT conditions, we have the following:
$\mathbf{D}^{*}\mathbf{Y}^{*}=\mathbf{D}^{*}\mathbf{I}\Rightarrow\operatorname{tr}(\mathbf{D}^{*}(\mathbf{Y}^{*}-\mathbf{I}))=0,$ (34)
$\mathbf{S}^{*}{\bm{\sigma}}^{*}=\mathbf{0}\Rightarrow\operatorname{tr}(\mathbf{S}^{*}\mathbf{Y}^{*})=0,$ (35)
where in eqn. (35), ${\bm{\sigma}}^{*}$ is the null space of $\mathbf{S}^{*}$. In order to ensure that $\mathbf{Y}^{*}$ is the unique solution, we require that $\lambda_{2}(\mathbf{S}^{*})>0$, i.e., the second smallest eigenvalue of $\mathbf{S}$. This comes from the rank-nullity theorem, i.e.,
$\text{rank}(\mathbf{S}^{*})+\text{null}(\mathbf{S}^{*})=n\Rightarrow\text{rank}(\mathbf{S}^{*})=n-1.$ (36)
To this end, we have the following:
$\operatorname{tr}({\mathbf{A}}\mathbf{Y})$ $\leq\mathcal{L}({\mathbf{A}},\mathbf{Y}^{*},\mathbf{S}^{*},\mathbf{D}^{*})$
$=\operatorname{tr}((\mathbf{S}^{*}-\mathbf{D}^{*}+{\mathbf{A}})\mathbf{Y})+\operatorname{tr}(\mathbf{D}^{*}\mathbf{I})$
$=\operatorname{tr}(\mathbf{D}^{*}\mathbf{I})=\operatorname{tr}(\mathbf{D}^{*}\mathbf{Y}^{*})$
$=\operatorname{tr}((\mathbf{S}^{*}+{\mathbf{A}})\mathbf{Y}^{*})=\operatorname{tr}({\mathbf{A}}\mathbf{Y}^{*}).$ (37)
Now, our goal is to prove that w.h.p. $\mathbf{S}^{*}\succcurlyeq 0$ with $\lambda_{2}(\mathbf{S}^{*})>0$. More specifically, we want to show that
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$ (38)
Before we proceed, we note that
$\mathds{E}[{\mathbf{A}}]$ $=(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$ (39)
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed with the following set of steps:
$\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}$ $=\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-\mathbf{x}^{T}\mathds{E}[{\mathbf{A}}]\mathbf{x}-\mathbf{x}^{T}({\mathbf{A}}-\mathds{E}[{\mathbf{A}}])\mathbf{x}$
$\geq\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-(1-2\zeta)p\mathbf{x}^{T}\mathbf{Y}^{*}\mathbf{x}+(1-2\zeta)p-\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|$
=$\overset{(a)}{=}\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}+(1-2\zeta)p-\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|$
$\geq\min_{i\in[n]}d_{i}^{*}+(1-2\zeta)p-\|{\mathbf{A}}-\mathds{E}[{\mathbf{A}}]\|$
$\overset{(b)}{\geq}\min_{i\in[n]}d_{i}^{*}+(1-2\zeta)p-c^{\prime}\sqrt{\log(n)},$ (40)
where step (a) holds from the unique optimality condition in (35). Note that $d_{i}$ is equal in distribution to $\sum_{i=1}^{n-1}X_{i}$, where $X_{i}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$. Step (b) follows that $\|\tilde{\mathbf{A}}-\mathds{E}[\tilde{\mathbf{A}}]\|\leq\tilde{c}^{\prime}\sqrt{\log(n)}$ with high probability, as convincingly demonstrated in [17]. The invocation of Lemma B.1 and the subsequent application of the union bound unequivocally validate the condition.
D.1 The Synergy of Combining Multiple Adjacency Matrices
In the case when we estimate $\bm{\sigma}$ via multiple observed adjacancy matrices, i.e., given ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{w}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, we have the MLE for $\bm{\sigma}$ resulted from SDP is
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\ \text{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)$
s.t. $\ \mathbf{Y}\succcurlyeq\mathbf{0},Y_{i,i}=1,\forall i\in[n].$
The Lagrangian function is written as follows:
$\mathcal{L}({\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{w},\mathbf{Y},\mathbf{S},\mathbf{D})=\operatorname{tr}\left(\sum_{t=1}^{w}{\mathbf{A}}_{t}\mathbf{Y}\right)+\operatorname{tr}(\mathbf{S}\mathbf{Y})-\operatorname{tr}(\mathbf{D}(\mathbf{Y}-\mathbf{I})),$ (41)
where $\mathbf{S}\succcurlyeq\mathbf{0}$ and $\mathbf{D}=\operatorname{diag}(d_{i})$. Then,
$\nabla_{\mathbf{Y}}\mathcal{L}=\sum_{t=1}^{w}{\mathbf{A}}_{t}+\mathbf{S}-\mathbf{D}=\mathbf{0}.$ (42)
In order to satisfy the first order stationary condition, we must satisfy the following (deterministic) condition:
$\mathbf{S}^{*}=\mathbf{D}^{*}-\sum_{t=1}^{w}{\mathbf{A}}_{t},$ (43)
Using Lemma 9 in [17], we have when ${\mathbf{S}}^{\ast}={\mathbf{D}}^{\ast}-\sum_{t=1}^{w}{\mathbf{A}}_{t}$ satisfies ${\mathbf{S}}^{\ast}\succcurlyeq\mathbf{0}$, $\lambda_{2}({\mathbf{S}}^{\ast})>0$ (second smallest eigenvalue) and
${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0,$
then $\hat{\mathbf{Y}}=\mathbf{Y}^{\ast}=\bm{\sigma}^{\ast}(\bm{\sigma}^{\ast})^{\top}$ is the unique solution. Therefore, it suffices to show ${\mathbf{S}}^{\ast}\succcurlyeq\mathbf{0}$ and $\lambda_{2}({\mathbf{S}}^{\ast})>0$. We let
${\mathbf{D}}^{\ast}=\text{diag}\{d_{i}^{\ast}\},\quad d_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{i}^{\ast}\sigma_{j}^{\ast}.$
By definition we have $d_{i}^{\ast}\sigma_{i}^{\ast}=\sum_{t=1}^{w}\sum_{j=1}^{n}{\mathbf{A}}_{t}(i,j)\sigma_{j}^{\ast}$, i.e., ${\mathbf{D}}^{\ast}\bm{\sigma}^{\ast}=(\sum_{t=1}^{w}{\mathbf{A}}_{t})\bm{\sigma}^{\ast}$, thus ${\mathbf{S}}^{\ast}\bm{\sigma}^{\ast}=0$. It remains to show that
$\operatorname{Pr}\left[\inf_{\mathbf{x}:\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}}\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}>0\right]\geq 1-o(1).$ (44)
Before we proceed, we note that
$\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]$ $=w(1-2\zeta)p(\mathbf{Y}^{*}-\mathbf{I}).$ (45)
Now, for any $\mathbf{x}$ such that $\|\mathbf{x}\|=1,\mathbf{x}\perp\bm{\sigma}^{*}$ (i.e., $\mathbf{x}^{T}\bm{\sigma}^{*}$ = 0), we proceed with the following set of steps:
$\mathbf{x}^{T}\mathbf{S}^{*}\mathbf{x}$ $=\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-\mathbf{x}^{T}\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\mathbf{x}-\mathbf{x}^{T}(\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}])\mathbf{x}$
$\geq\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}-w(1-2\zeta)p\mathbf{x}^{T}\mathbf{Y}^{*}\mathbf{x}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|$
=$\overset{(a)}{=}\mathbf{x}^{T}\mathbf{D}^{*}\mathbf{x}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|$
$\geq\min_{i\in[n]}d_{i}^{*}+w(1-2\zeta)p-\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|,$ (46)
where step (a) holds since $\mathbf{x}\perp\bm{\sigma}^{*}$. Note that each $d_{i}^{\ast}$ is equal in distribution to $\sum_{j=1}^{w(n-1)}X_{j}$, where $X_{j}\overset{i.i.d.}{\sim}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0}$.
First, we have for $m=w(n-1)$, and $k_{n}>0$,
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg{)}\leq\exp(-m\ell(k_{n}/m)),$
where $\ell(x)=\sup_{t\geq 0}-tx-\log\mathds{E}\left[e^{-tX_{1}}\right]$.
The term $\mathds{E}\left[e^{-tX_{1}}\right]$ is obtained as
$\mathds{E}\left[e^{-tX_{1}}\right]=1+p\left[e^{-t}(1-{\zeta})+e^{t}{\zeta}-1\right].$
Note that the above equation is a concave function of the parameter $t$. The optimum $t^{*}$ that maximizes (52) is given as follows:
$-x+\frac{{p}\left[e^{-t^{*}}(1-{\zeta})-e^{t^{*}}{\zeta}\right]}{1+{p}\left[e^{-t^{*}}(1-{\zeta})+e^{t^{*}}{\zeta}-1\right]}=0.$ (47)
For $x=o(1)$, the optimal parameter $t^{*}$ for the Chernoff bound is obtained as
$t^{*}$ $=\frac{1}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(1).$ (48)
Plugging the value of $t^{*}$, it gives us the following:
$\ell(k_{n}/m)$ $=-t^{*}\frac{k_{n}}{m}-\log\mathds{E}\left[e^{-t^{*}X_{1}}\right]$
$=-\frac{k_{n}}{2m}\log\frac{1-{\zeta}}{{\zeta}}+a\bigg{(}\sqrt{1-{\zeta}}-\sqrt{{\zeta}}\bigg{)}^{2}\frac{\log n}{n}+o(k_{n}/m)$
where the last equality follows due to the Taylor expansion of $\log(1-x)$ around $x=0$, and $p=a\log(n)/n$.
Thereby
$\exp(-m\ell(k_{n}/m))=\exp(-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}\log n+\frac{k_{n}}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(k_{n}))$
choose $k_{n}=(1+o(1))\frac{\log n}{\log\log n}$ yields $\exp(-m\ell(k_{n}/m))=\exp(-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}\log n+\frac{k_{n}}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(k_{n}))=n^{-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$. Applying the union bound on $i$, we have with probability $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$,
$\min_{i}d_{i}^{\ast}\geq\frac{\log n}{\log\log n}.$
Second, we have
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq\sum_{t=1}^{w}\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|$
by Theorem 9 in [17] we have $\operatorname{Pr}\{\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|\leq c^{\prime}\sqrt{a\log n}\}\geq 1-n^{c}$, we have with probability at least $1-wn^{-c}$,
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq wc^{\prime}\sqrt{a\log n},$
for $p=a\frac{\log n}{n}$.
Combing these together we have with probability at least $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}-wn^{-c}$, we have (46) $\geq\frac{\log n}{\log\log n}+w(1-2\zeta)p-wc^{\prime}\sqrt{a\log n}>0$ for $n$ sufficiently large. And therefore the model satisfies the exact recovery condition.
Appendix E Analysis of the Graph Perturbation Mechanism
In this section, we provide analysis for the Ternary Randomized Resonse Mechanism used in Section 4.1. In this mechanism, when $A_{i,j}=x$ where $x\in\{0,+1,-1\}$, the mechanism outputs $\tilde{A}_{i,j}=y$, where $y\in\{0,+1,-1\}$ with probability $q_{x}(y)\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=y|{\mathbf{A}}_{i,j}=x)$. We set $q_{x}(y)=c_{2}$, $\forall x\neq y$, and $q_{x}(y)=c_{1}=1-2c_{2}$, $\forall x=y$. We have the following:
$\tilde{p}_{+1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=+1|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{1}p(1-\zeta)+c_{2}p\zeta+c_{2}(1-p),$
$\tilde{p}_{-1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=-1|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{1}p\zeta+c_{2}(1-p),$
$\tilde{p}_{0}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=0|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p),$
$\tilde{q}_{+1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=+1|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{1}p\zeta+c_{2}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}_{-1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=-1|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{2}p\zeta+c_{1}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}_{0}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=0|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p).$
Remark E.1 (CBM model after ternary-randomized perturbation).
We note that after the above ternary-randomized mechanism, the distribution of $\tilde{{\mathbf{A}}}$ is still a CBM model. More specifically, for the original CBM model ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, the perturbed $\tilde{\mathbf{A}}$ follows the CBM model ${\mathsf{CBM}}(\bm{\sigma},\tilde{p},\tilde{\zeta})$ with the same community label vector $\bm{\sigma}$ and new parameters
$\tilde{p}=1-c_{1}(1-p)-c_{2}p=2c_{2}+p(1-3c_{2}),\,\tilde{\zeta}=\frac{c_{2}(1-p\zeta)+c_{1}p\zeta}{1-c_{1}(1-p)-c_{2}p}=\frac{c_{2}+p\zeta(1-3c_{2})}{2c_{2}+p(1-3c_{2})}.$ (49)
Remark E.2 (KL Divergence after Perturbation).
We first analyze the change in the KL divergence between the pre- and post-change distributions after the perturbation. The KL divergence for the raw data distribution (before perturbation) can be derived as (see Remark 5.2 for details),
$I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}\right].$
After the perturbation, we have similarly
$\tilde{I}(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\left(\frac{1}{2}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}\right)\tilde{p}(1-2\tilde{\zeta})\left[{n\choose 2}-\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}\right].$
Substitute the parameter $\tilde{p}$ and $\tilde{\zeta}$ in (49), and using the setting that $c_{1}=1-2c_{2}$, we have
$\left(\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}\right)\tilde{p}(1-2\tilde{\zeta})=\left(\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}\right)\cdot p(1-2\zeta)(1-3c_{2}).$ (50)
We note that $\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}$ is a positive and decreasing function of $c_{2}$ when $\zeta<\frac{1}{2}$ and a negative and increasing function when $\zeta>\frac{1}{2}$, the same is the function $p(1-2\zeta)(1-3c_{2})$. Moreover, when $c_{2}=0$ we have the KL divergence remaining unchanged. Combining these together, we have (50) is a nonnegative function and it is a decreasing function of $c_{2}$, with the largest value obtained when $c_{2}=0$ and the largest value equal to $I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})$. This demonstrates the tradeoff between detection and privacy. This being said, for a given privacy constraint, we may select the smallest value of $c_{2}$ value that satisfies this constraint; such choice ensures the best detection performance after the perturbation.
Note that, under $\epsilon$-edge LDP, we require $c_{2}/c_{1}\in[e^{-\epsilon},e^{\epsilon}]$, which yields the smallest possible value for $c_{2}$ being $c_{2}=\frac{1}{e^{\epsilon}+2}$, which is the value we choose in Section 4.1.
Next we derive the condition for exact recovery under the privacy mechanism.
Lemma E.1.
(Chernoff Bound for the Perturbed Graph via Randomized Response) Let $\{X_{i}\}_{i=1}^{m}$ be a collection of independent random variables, where $m=n+o(n)$ and each $X_{i}\overset{i.i.d.}{\sim}\tilde{p}_{+1}\delta_{+1}+\tilde{p}_{-1}\delta_{-1}+\tilde{p}_{0}\delta_{0}$. Then,
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg{)}\leq\exp(-m\ell(k_{n}/m)),$ (51)
where $k_{n}=(1+o(1))\frac{\log(n)}{\log(\log(n))}$ and $\ell(k_{n}/m)=\sup_{t\geq 0}-tx-\log\mathds{E}\left[e^{-tX_{1}}\right]$.
The term $\mathds{E}\left[e^{-tX_{1}}\right]$ is obtained as
$\mathds{E}\left[e^{-tX_{1}}\right]$ $=1+\tilde{p}\left[e^{-t}(1-\tilde{\zeta})+e^{t}\tilde{\zeta}-1\right].$ (52)
Note that the above equation is a concave function of the parameter $t$. The optimum $t^{*}$ that maximizes (52) is given as follows:
$-x+\frac{\tilde{p}\left[e^{-t^{*}}(1-\tilde{\zeta})-e^{t^{*}}\tilde{\zeta}\right]}{1+\tilde{p}\left[e^{-t^{*}}(1-\tilde{\zeta})+e^{t^{*}}\tilde{\zeta}-1\right]}=0.$
For $x={k_{n}}/{m}$, the optimal parameter $t^{*}$ for the Chernoff bound is obtained as
$t^{*}$ $=\frac{1}{2}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}+o(1).$
Plugging the value of $t^{*}$, it gives us the following:
$\ell(k_{n}/m)$ $=-t^{*}\frac{k_{n}}{m}-\log\mathds{E}\left[e^{-t^{*}X_{1}}\right]$
$=-\frac{k_{n}}{2m}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}-\log\left[1-\tilde{p}\bigg{(}\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg{)}^{2}\right]+o(k_{n}/m)$
=$\overset{(a)}{=}\tilde{p}\bigg{(}\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg{)}^{2}+o(\log(n)/n)$
$=\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 43.36243pt\times\bigg{(}\sqrt{\frac{1+(1-\zeta)p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\sqrt{\frac{1+\zeta p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\bigg{(}\sqrt{\frac{(1-\zeta)p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\sqrt{\frac{\zeta p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\frac{1}{\sqrt{2+p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\frac{p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}-\frac{1}{\sqrt{p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\frac{p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\overset{(b)}{\geq}\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\alpha\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}\bigg{)}^{2}\log(n)/n+o(\log(n)/n),$ (53)
where step $(a)$ follows due to the Taylor expansion of $\log(1-x)$ around $x=0$, and $p=a\log(n)/n$ where the first term in step (a) behaves as $o(\log(n)/n)$. Step (b) holds true whenever $a>\frac{2(n^{3/2}-n)}{(n^{c}-1)\log(n)}$, where $\alpha=1-\frac{1}{\sqrt{n}}$, $\epsilon=c\log(n)$ and $c>0$. After applying the union bound, we get the following sufficient conditions for exact recovery:
$a\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}>\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\times\left[1/\alpha-\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\times\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}\right].$
From the above condition, we conclude that the privacy leakage $\epsilon$ should behave as $\Omega(\log(n))$. A more stringent condition yields
$a\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}>\frac{1}{\alpha}\times\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right).$
Remark E.3.
In conclusion, the analysis of the lower bounds on the parameter $a$ for different formulations of $\alpha$, specifically $\alpha=1-\frac{1}{\sqrt{n}}$ and $\alpha=1-\frac{1}{n}$, reveals significant variations in the required values of $a$ for a given $c$. Notably, for a fixed $c$, the choice of $\alpha=1-\frac{1}{\sqrt{n}}$ generally results in a tighter (or lower) lower bound for $a$ across a range of $n$ values. This tighter bound is preferable as it allows for more flexibility and potentially less stringent requirements on the parameter $a$ while maintaining the desired privacy level defined by $\epsilon=c\log(n)$. Thus, in scenarios where achieving an optimal balance between privacy and parameter constraints is crucial, opting for $\alpha=1-\frac{1}{\sqrt{n}}$ would be a more strategic choice.
Figure 4: Impact of the Privacy Scaling Coefficient $c$ on the Required Lower Bound of $a$: This plot illustrates the relationship between the lower bound of the parameter $a$ and the number of samples $n$, under the privacy condition $\alpha=1-\frac{1}{\sqrt{n}}$ and the privacy budget $\epsilon=c\log(n)$, where $c>0$. The analysis reveals that an increase in the value of $c$ leads to a reduction in the minimum required value of $a$, indicating a trade-off between the privacy scaling coefficient and the parameter $a$.
Appendix F Fundamental Limits of Private Community Recovery
The proof technique is inspired by the packing argument [33] in which we provide a lower bound for estimation error under $\epsilon$-edge DP.
Lemma F.1.
(Packing lower bound for graph datasets) Let $\mathbf{A}_{1},\mathbf{A}_{2},\cdots\mathbf{A}_{m}\in\{0,-1,1\}^{n\times n}$ be a set of $m$ graph datasets . For any pair of datasets $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$, we have $\operatorname{Ham}(\mathbf{A}_{i},\mathbf{A}_{j})\leq d$. Let $\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{m}$ be a set of $m$ disjoint subsets. If there exists an $\epsilon$-DP mechanism $\mathcal{M}:\mathcal{G}\rightarrow\mathcal{S}$ such that $\operatorname{Pr}(\mathcal{M}(G_{i})\in\mathcal{S}_{i})\geq 9/10$ for every $i\in[m]$, then $\epsilon=\Omega(\log(m)/d)$.
Remark F.1.
Note that $d={n\choose 2}$ is the worst case upper bound on the Hamming distance between any two graphs (i.e., adjacency matrices). We next derive similar type of lower bounds that are stronger by replacing the worst-case distance with the expected hamming distance between any two random graphs generated from a joint distribution.
F.1 Converse for $\epsilon$-edge DP Community Recovery Algorithms
We establish a rigorous lower bound for all differentially private community recovery algorithms operating on graphs generated from the Constrained Block Models (CBMs). Our methodology closely follows the frameworks outlined in [33, 6], focusing on the notion of edge DP. Furthermore, we precisely define the classification error rate as
${\operatorname{err}}\operatorname{rate}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})$ $=\frac{1}{n}\times\min\{\operatorname{Ham}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}}),\operatorname{Ham}(-\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})\}.$ (54)
Let us consider a series of pairwise disjoint sets $\mathcal{S}_{i}$ for $i=1,2,\ldots,m$. Each set $\mathcal{S}_{i}$ contains vectors $\bm{u}\in\{\pm 1\}^{n}$, where $n$ is the vector dimension. A vector $\bm{u}$ is included in $\mathcal{S}_{i}$ if the error rate $\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma})$ with a fixed vector $\bm{\sigma}$ does not exceed the threshold $\beta$. This is formally expressed as:
$\mathcal{S}_{i}=\{\bm{u}\in\{\pm 1\}^{n}:\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma})\leq\beta\},\quad i=1,2,\ldots,m,$ (55)
where $\mathcal{S}_{i}$’s are pairwise disjoint sets.
Further, for each index $i$, let $P_{i}$ be the distribution over $n$-vertex graphs generated by an CBM with parameters $\zeta$, $p$ and $\bm{\sigma}^{i}$, in short $\operatorname{CBM}(\zeta,p,\sigma^{i})$. We next derive the necessary conditions for
${\operatorname{Pr}}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})\geq 1-\eta$ (56)
as a function of the CBM parameters and the privacy budget. Note that the randomness here is taken over the randomness of graph $G$ that is generated from $\operatorname{CBM}(\zeta,p,\sigma^{i})$. Without loss of generality, let us consider a graph $\mathbf{A}\sim\operatorname{CBM}(\zeta,p,\sigma^{1})$ that is generated from the ground truth labeling vector $\bm{\sigma}^{*}=\bm{\sigma}^{1}$. Further, for this case, we want to derive the necessary conditions for any $\epsilon$-edge DP recovery algorithms that
${\operatorname{Pr}}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{1})\geq 1-\eta\Rightarrow\sum_{i=2}^{m}{\operatorname{Pr}}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})\leq\eta,$ (57)
where $\mathbf{A}\sim\operatorname{CBM}(\zeta,p,\sigma^{1})$.
We next individually lower bound each term in eqn. (57). To do so, we first invoke the group privacy property of DP [12] and show that for any two adjacency matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$, we have
$\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S})\leq e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}),$ (58)
for any measurable set $\mathcal{S}\subseteq\{\pm 1\}^{n}$. For each $i=1,2,\cdots,m$, taking the expectation with respect to the coupling distribution $\Pi(\mathbf{A},\mathbf{A}^{\prime})$ between $\mathbf{A}$ and $\mathbf{A}^{\prime}$ and setting $\mathcal{S}=\mathcal{S}_{i}$, yields the following:
$\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})\right]$ $\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
$\Rightarrow\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})$ $\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
$\overset{(a)}{\leq}\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}$
×$\hskip 93.95122pt\times\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}^{2}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]\right)^{1/2}$
$\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow 1-\eta$ $\overset{(b)}{\leq}\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow(1-\eta)^{2}$ $\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)\times\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i}),$ (59)
where step (a) follows from applying Cauchy-Schwartz inequality. In step (b), we invoked the condition in eqn. (56).
F.2 Characterizing the randomness of the Hamming distance between two random adjacency matrices
We first note that the Hamming distance between two labeling vectors $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ (each of size $n$) directly determines how many rows in the adjacency matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$ are generated from the same versus different distributions. More precisely, we have two cases: case $(1)$: $\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})$ rows in $\mathbf{A}$ and $\mathbf{A}^{\prime}$ have elements generated from different distributions, and case $(2)$: $n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})$ rows have elements from the same distribution.
Case $(1)$: In this case, the probability the corresponding elements in the two matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$ are different is $\bar{q}=1-((1-\zeta)\zeta p^{2}+(1-\zeta)\zeta p^{2}+(1-p)^{2})$.
Case $(2)$: In this case, the probability the corresponding elements in the two matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$ are same is $\bar{p}=1-((1-\zeta)^{2}p^{2}+\zeta^{2}p^{2}+(1-p)^{2})$.
We next focus in calculating the term $M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)\triangleq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]$ with the following set of steps:
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$ $=\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))}\times\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{different dist.}}(2\epsilon)\right)^{\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})},$ (60)
where,
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{samedist.}}(2\epsilon)$ $=e^{2\epsilon}\bar{p}+(1-\bar{p}),$ (61)
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{different dist.}}(2\epsilon)$ $=e^{2\epsilon}\bar{q}+(1-\bar{q}).$ (62)
We then can readily show that,
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$ $\leq\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})}$
$=\left(-2p^{2}\zeta(\zeta-1)+(p-1)^{2}+\left(2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1\right)e^{2\epsilon}\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})}$
$=(1-p^{\prime}+p^{\prime}e^{2\epsilon})^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})},$ (63)
where $p^{\prime}=2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1$. Plugging (63) in (59) yields the following:
$\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A}^{\prime})\in\mathcal{S}_{i})$ $\geq\frac{(1-\eta)^{2}}{M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)}.$ (64)
F.3 Lower bounding the packing number $m$ with respect to $\operatorname{err}\operatorname{rate}$:
Building upon the framework established in [6], we can readily demonstrate that
$m\geq\frac{1}{2}\cdot\frac{|\mathcal{B}_{\operatorname{Ham}}(\bm{\sigma}^{*},4\beta n)|}{|\mathcal{B}_{\operatorname{Ham}}(\bm{\sigma}^{*},2\beta n)|},$ (65)
where $\mathcal{B}_{\operatorname{Ham}}(\bm{\sigma}^{*},t\beta n)=\{\bm{\sigma}\in\{\pm 1\}^{n}:\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{*})\leq\beta\}$ and $|\mathcal{B}_{\operatorname{Ham}}(\bm{\sigma}^{*},t\beta n)|$ is the number of vectors within the Hamming distance of $t\beta n$ from $\bm{\sigma}^{*}$ for $t>0$. It includes all vectors that can be obtained by flipping any $t\beta n$ elements of $\bm{\sigma}^{*}$. Thus, we can further lower bound $m$ as
$m\geq\frac{1}{2}\cdot\frac{{n\choose 4\beta n}}{{n\choose 2\beta n}}\geq\frac{1}{2}\cdot\left(\frac{1}{8e\beta}\right)^{2\beta n}.$ (66)
Recall that, we have
$(m-1)\cdot\frac{(1-\eta)^{2}}{M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)}\leq\eta.$ (67)
Taking the logarithm for both sides of (67), it yields
$8\beta n(n-8\beta n)\log(1-p^{\prime}+p^{\prime}e^{2\epsilon})\geq 2\beta n\log\left(\frac{1}{8e\beta}\right)+\log\left(\frac{1}{\eta}\right)$
$\Rightarrow$ $\log(1-p^{\prime}+p^{\prime}e^{2\epsilon})\geq\frac{\log\left(\frac{1}{8e\beta}\right)}{4(n-8\beta n)}+\frac{\log\left(\frac{1}{\eta}\right)}{8\beta n(n-8\beta n)}$
$\Rightarrow$ $\left(e^{2\epsilon}-1\right)p^{\prime}\geq\frac{\log\left(\frac{1}{8e\beta}\right)}{4(n-8\beta n)}+\frac{\log\left(\frac{1}{\eta}\right)}{8\beta n(n-8\beta n)}$
$\overset{(a)}{\Rightarrow}$ $\left(e^{2\epsilon}-1\right)p^{\prime}\geq\frac{(\Delta+1)\log(n)-\log(8e)}{4(n-8)}$
$\Rightarrow$ $\epsilon\geq\frac{1}{2}\log\left[1+\frac{(\Delta+1)\log(n)-\log(8e)}{p^{\prime}(4n-32)}\right],$ (68)
where $p^{\prime}=2a\times\frac{\log(n)}{n}+a^{2}(2\zeta^{2}-2\zeta-1)\times\frac{\log^{2}(n)}{n^{2}}=2a\frac{\log(n)}{n}+\mathcal{O}\left(\frac{\log^{2}(n)}{n^{2}}\right)$ In step (a), we set $\beta=n^{-1}$ and $\eta=n^{-\Delta}$, where $\Delta>0$.
Appendix G Performance of the Private Online Change Detection Algorithm
In this section, we provide a detailed upper bound for the detection delay for the detection algorithm in Section 4. For simplicity, we perform the analysis on the data sequence $\{{\mathbf{A}}_{t},t\in\mathbb{N}\}$, and here ${\mathbf{A}}_{t}$ can stands for both the original data and the perturbed private data, since they both satisfy the CBM model. Denote the KL information number $I_{0}={\mathbb{E}}_{0}^{\bm{\sigma}^{\text{post}}}\left[\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}})\right]$
Average Run Length
We begin by considering the average run length, which characterizes average false alarm period when there is no change in the community structure.
Proof of Lemma 5.2.
We compute the average false alarm period using similar ideas as in Lemma 8.2.1 in [32]. For $t>w$ let us define a Shiryaev-Roberts like statistic $\{R_{t}\}$ as
$R_{t}=(R_{t-1}+1)e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},~{}R_{w}=0.$
It is important to note that $\{R_{t}-t\}$ is a martingale with respect to the ${\mathbb{E}}_{\infty}$ measure since
${\mathbb{E}}_{\infty}[R_{t}-t|{\mathcal{F}}_{t-1}]={\mathbb{E}}_{\infty}\left[(R_{t-1}+1)e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})}-t\Big{|}{\mathcal{F}}_{t-1}\right]=R_{t-1}-(t-1).$
The last equality is true because given ${\mathcal{F}}_{t-1}$ we have $\bm{\hat{\sigma}}_{t-1}$ fixed and is independent of the observation ${\mathbf{A}}_{t}$. Since $\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})$ is a valid log density ratio for ${\mathbf{A}}_{t}$, we have the its expectation under ${\mathbb{E}}_{\infty}$ regime equals to one. The martingale property of $\{R_{t}-t\}$ and usage of Optional Sampling allows us to write for any stopping time $T$ with finite expectation that
${\mathbb{E}}_{\infty}[R_{T}-T]={\mathbb{E}}_{\infty}[R_{w}-w]=-w~{}\Rightarrow~{}{\mathbb{E}}_{\infty}[T]-w={\mathbb{E}}_{\infty}[R_{T}].$ (69)
Recall that the detection statistics updated in (14), after exponentiation, can be equivalently written as
$e^{S_{t}}=\max\{e^{S_{t-1}},1\}e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},~{}e^{S_{w}}=1.$
Using induction, the fact that $e^{S_{w+1}}=R_{w+1}$ and that for $x\geq 0$, $x+1\geq\max\{x,1\}$, it is straightforward to prove that for $t>w$ we have $R_{t}\geq e^{S_{t}}$. With this observation and using (69) we can now write
$e^{b}\leq{\mathbb{E}}_{\infty}[e^{S_{T}}]\leq{\mathbb{E}}_{\infty}[R_{T}]={\mathbb{E}}_{\infty}[T]-w\leq{\mathbb{E}}_{\infty}[T],$
which proves the desired inequality. ∎
Given the above Lemma, it is obvious to see that we can choose the detection threshold
$b=\log\gamma$
such that the average run length constraint ${\mathbb{E}}_{\infty}[T]\geq\gamma$ is satisfied. Therefore, we adopt this threshold in subsequent delay analysis.
Detection Delay
In the following, we provide theoretical guarantees on the detection delay of the procedure (14). We first define the quantities, $\hat{I}_{0}={\mathbb{E}}_{0}^{\bm{\sigma}^{\text{post}}}\left[\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})\right]$, which resemble the KL information number $I_{0}$ above.
Theorem G.1 (WADD Upper Bound).
Assume $\hat{I}_{0}>0$. Under the assumptions of Theorem 5.4, we have the following upper bound for the worst-case performance of the detection procedure (17),
${\mathbb{E}}_{0}^{\bm{\theta}^{\text{post}}}[T]\leq\frac{\log\gamma}{\hat{I}_{0}}(1+o(1)).$ (70)
Proof.
The proof follows from Theorem 1 of [36], after observing that $\log\gamma$ dominates the numerator term under our assumption. ∎
In the following, we analyze the empirical quantities $\hat{I}_{0}$ and establish the asymptotic optimality result. The key component in the analysis is the estimation error of the MLE $\hat{\bm{\sigma}}_{t}$. We first establish the probability of false estimation.
Lemma G.1 (Estimating $\bm{\sigma}$).
Under the post-change regime, if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$, the estimate $\bm{\hat{\sigma}}$ resulted from $w$ i.i.d. post-change samples (adjacency matrices) satisfies, as $n\to\infty$,
$\operatorname{Pr}(\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}})\geq 1-wn^{-\Omega(1)}.$
Proof.
From the analysis in Section D, we write $\bar{\mathbf{A}}=\sum_{i=1}^{w}{\mathbf{A}}_{i}$. Following the result of Theorem $9$ in [17], for any constant $c>0$, there exists $c^{\prime}>0$ such that for any $n\geq 1$,
$\Pr\left\{\left\|\bar{\mathbf{A}}-{\mathbb{E}}[\bar{\mathbf{A}}]\right\|_{2}\leq c^{\prime}w\sqrt{np}\right\}\geq 1-wn^{-c}.$
Following the proof to Theorem 6 in [17], we have the desired result. ∎
Under Lemma G.1, we have
$\hat{I}_{0}=I_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}$
note that $\mathbb{E}[\hat{I}_{0}]=\mathbb{E}[\log\frac{p(A_{t};\hat{\bm{\sigma}}_{t})}{p(A_{t};\bm{\sigma}^{\text{pre}})}]=\mathbb{E}[\log\frac{p(A_{t};\hat{\bm{\sigma}}_{t})}{p(A_{t};\bm{\sigma}^{\text{post}})}+\log\frac{p(A_{t};\bm{\sigma}^{\text{post}})}{p(A_{t};\bm{\sigma}^{\text{pre}})}]=I_{0}-I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})$.
We note that when estimating the post-change community structure within the support that satisfying Assumption 1, we have $I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})\leq 2\Delta n\left(\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)$ (see Lemma G.2 below, together with Remark 5.2), and thus $|I_{0}-I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})|\leq 3\Delta n\left(\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)$. Note that
$\frac{\hat{I}_{0}}{I_{0}}=\frac{I_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}}{I_{0}}=1-O(wn^{-c}),$
and when combined with (70), we obtain the desired first-order optimality in Theorem 5.4.
Lemma G.2.
Under Assumption 1, we have
(${n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\leq 2\Delta n,$ (71)
where $C_{\bm{\sigma},\bm{\sigma}^{\prime}}:=\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma^{\prime}}_{i}\bm{\sigma^{\prime}}_{j}$.
Indeed, when there are $\Delta$ users who change their membership between $+1$ and $-1$ after the change, we have the quantity (${n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}$ is upper bounded by $2\Delta(n-\Delta)$.
Appendix H Fundamental Limits of Private Online Change Detection
Recall that we denote ${\mathbf{A}}$ as the original adjacency matrix, and $\tilde{\mathbf{A}}$ as the perturbed matrix yielded by a $\epsilon$ edge-LDP randomized mechanism. We denote $\tilde{p}_{\infty}$ as the distribution for the perturbed matrix $\tilde{\mathbf{A}}$ in the pre-change regime, and let $\tilde{p}_{0}$ be the distribution in the post-change regime.
In our setting when only the community structure changes, while $p,\zeta$ remain unchanged and known, the KL equals to
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})=\sum_{i<j:\ \sigma^{\text{post}}_{i}\sigma^{\text{post}}_{j}\neq\sigma^{\text{pre}}_{i}\sigma^{\text{pre}}_{j}}\sum_{z}\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}.$
We note that for pair $i<j$ such that $\sigma^{\text{post}}_{i}=\sigma^{\text{post}}_{j},\sigma^{\text{pre}}_{i}\neq\sigma^{\text{pre}}_{j}$, we have
$\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)=q_{1}(z)p(1-\zeta)+q_{-1}(z)p\zeta+q_{0}(z)(1-p),$
and
$\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)=q_{1}(z)p\zeta+q_{-1}(z)p(1-\zeta)+q_{0}(z)(1-p).$
Lemma H.1.
For the perturbed data distribution $\tilde{p}_{1}$ and $\tilde{p}_{0}$, we have for each $z\in\{+1,-1,0\}$,
$\left|\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}\right|\leq c_{\epsilon}(e^{\epsilon}-1)p(1-2\zeta),$
where $c_{\epsilon}=\min\{2,e^{\epsilon}\}$.
Proof.
The proof follows from Lemma 1 of [11]. We first have $|\log\frac{a}{b}|\leq\frac{|a-b|}{\min\{a,b\}}$ for $a,b\in\mathbb{R}_{+}$. Indeed, for any $x>0$, $\log(x)\leq x-1$. Set $x=a/b$ and $x=b/a$ yields
$\log\frac{a}{b}\leq\frac{a}{b}-1=\frac{a-b}{b},\quad\log\frac{b}{a}\leq\frac{b}{a}-1=\frac{b-a}{a}.$
Using the first inequality for $a>b$ and the second for $a<b$ yields $|\log\frac{a}{b}|\leq\frac{|a-b|}{\min\{a,b\}}$.
Let $a=\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)$ and $b=\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)$, we have
$\left|\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}\right|\leq\frac{|\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)-\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)|}{\min\{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z),\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)\}}.$
We further have
$|\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)-\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)|=|q_{1}(z)-q_{-1}(z)|\cdot|p(1-2\zeta)|.$
For any $x\in\{-1,+1,0\}$ we have
$|q_{1}(z)-q_{-1}(z)|=|q_{1}(z)-q_{x}(z)+q_{x}(z)-q_{-1}(z)|\leq|q_{1}(z)-q_{x}(z)|+|q_{x}(z)-q_{-1}(z)|.$
This means we have
$|q_{1}(z)-q_{-1}(z)|\leq 2\min_{x}q_{x}(z)\max_{x^{\prime}}\left|\frac{q_{x^{\prime}}(z)}{q_{x}(z)}-1\right|.$
Similarly we have
$|q_{1}(z)-q_{-1}(z)|\leq q_{-1}(z)\left|\frac{q_{1}(z)}{q_{-1}(z)}-1\right|\leq e^{\epsilon}\min_{x}q_{x}(z)\left|\frac{q_{1}(z)}{q_{-1}(z)}-1\right|.$
Combing these together, and using the result that $e^{\epsilon}-1\geq 1-e^{-\epsilon}$, we have
$|q_{1}(z)-q_{-1}(z)|\leq c_{\epsilon}\min_{x}q_{x}(z)(e^{\epsilon}-1).$
Therefore we have
$\left|\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}\right|$ $\leq\frac{|\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)-\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)|}{\min\{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z),\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)\}}$
$\leq\frac{c_{\epsilon}\min_{x}q_{x}(z)(e^{\epsilon}-1)\cdot|p(1-2\zeta)|}{\min_{x}q_{x}(z)}$
$=c_{\epsilon}(e^{\epsilon}-1)\cdot|p(1-2\zeta)|.$
Proof of Lemma 5.3.
From the above lemma, we have
$\sum_{z}\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}$ $\leq\sum_{z}(\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)-\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z))\log\frac{\tilde{p}_{0}(\tilde{\mathbf{A}}_{i,j}=z)}{\tilde{p}_{\infty}(\tilde{\mathbf{A}}_{i,j}=z)}$
$\leq c_{\epsilon}^{2}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}\sum_{z}\min_{x}q_{x}(z)\leq c_{\epsilon}^{2}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}.$
The same inequality hold for pair $i<j$ such that $\sigma^{\text{post}}_{i}\neq\sigma^{\text{post}}_{j},\sigma^{\text{pre}}_{i}=\sigma^{\text{pre}}_{j}$. Therefore we have
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})\leq c_{\epsilon}^{2}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}\left[{n\choose 2}-\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}\right].$